### v1

# Measurement induced phase transition in a quantum Ising model by periodic measurements

Abstract
Measurement-induced phase transitions are often studied in random quantum circuits, with local measurements performed with a certain probability. We present here a model where a global measurement is performed with certainty at every time-step of the measurement protocol. Each time step, therefore, consists of evolution under the transverse Ising Hamiltonian for a time $\tau$, followed by a measurement that provides a “yes/no” answer to the question, “Are all spins up?”. The survival probability after $n$ time-steps is defined as the probability that the answer is “no” in all the $n$ time-steps. For various $\tau$ values, we compute the survival probability, entanglement in bipartition, and the generalized geometric measure, a genuine multiparty entanglement, for a chain of size $L\sim 26$, and identify a transition at $\tau_{c}\sim 0.2$ for field strength $h=1/2$. We then analytically derive a recursion relation that enables us to calculate the survival probability for system sizes up to 1000, which provides evidence of a scaling $\tau_{c}\sim 1/\sqrt{L}$. The transition at finite $\tau_{c}$ for $L\sim 28$ seems therefore to recede to $\tau_{c}=0$ in the thermodynamic limit. Additionally, at large time-steps, survival probability decays logarithmically only when the ground state of the Hamiltonian is paramagnetic. Such decay is not present when the ground state is ferromagnetic.
I Introduction
A paradigmatic model for studying measurement-induced phase transition (MIPT) is a quantum gate circuit of some suitable structure in which measurements are performed locally Li et al. (2018); Skinner et al. (2019); Rakovszky et al. (2019); Li et al. (2019); Chan et al. (2019); Szyniszewski et al. (2019); Bao et al. (2020); Choi et al. (2020); Gullans and Huse (2020a, b); Jian et al. (2020); Zabalo et al. (2020); Iaconis et al. (2020); Turkeshi et al. (2020); Zhang et al. (2020); Szyniszewski et al. (2020); Nahum et al. (2021); Ippoliti et al. (2021); Ippoliti and Khemani (2021); Lavasani et al. (2021a, b); Sang and Hsieh (2021); Block et al. (2022); Sharma et al. (2022); Agrawal et al. (2022); Barratt et al. (2022); Jian et al. (2023); Kelly et al. (2023), (for review, see Fisher et al. (2023)). The time evolution of the system consists of unitary evolution under some operator $U$ for an interval of time, followed by (projective or weak) measurement of some suitable local operator with probability ${\mathcal{P}}$. After every time step, the bipartite entanglement entropy $S(t)$ is measured. It is found that in the asymptotic limit of large size and long time, for small values of ${\mathcal{P}}$, $S(t)$ increases linearly with time, while for large values of ${\mathcal{P}}$ it saturates after some time. At some intermediate value ${\mathcal{P}}={\mathcal{P}}_{c}$, the entropy increases logarithmically with time, and this is considered to be a signature of a phase transition that has been induced by the process of measurement. The occurrence of such transitions has been confirmed in numerous systems in the last 5 years, and most of them are ‘small’ in size ($<40$) Skinner et al. (2019); Chan et al. (2019); Szyniszewski et al. (2019); Bao et al. (2020); Choi et al. (2020), while some are on larger sizes ($>100$) Li et al. (2018, 2019); Gullans and Huse (2020a, b). The location of the transition point ${\mathcal{P}}_{c}$ is believed to be independent of size. Apart from the random circuit analysis, this type of transition is also captured in the transverse field Ising model  Boorman et al. (2022); Sierant et al. (2022a); Malakar et al. (2024); Su et al. (2024); Tirrito et al. (2023); Roser et al. (2023); Biella and Schiró (2021); Banerjee and Sengupta (2024); Paviglianiti and Silva (2023), fermionic systems Buchhold et al. (2021); Poboiko et al. (2024); Müller et al. (2022); Chatterjee and Modak (2024); Minato et al. (2022); Jin and Martin (2024); Jian et al. (2021), many-body localized systems Lunt and Pal (2020) and higher-dimensional systems Turkeshi et al. (2020); Lunt et al. (2021); Nahum et al. (2021); Sierant et al. (2022b); Feng et al. (2023). This volume-to-area law transition is also observed in experiments with trapped ions and superconducting qubits Noel et al. (2022); Koh et al. (2023); AI and Collaborators (2023); Kamakari et al. (2024). Dependence of MIPT on the entangling power of unitary gate has been explored in Manna et al. (2024). Two essential ingredients of these studies are the local nature of the measurement and the randomness in whether the measurement will be performed or not at a particular time step. It would therefore be interesting to investigate what happens when these two features are removed. More importantly, for an integrable system, if one can go to a larger size ($\sim 1000$), it would be useful to check if the transition point depends on system size.
In this work, we shall consider non-random global projective measurement performed on a quantum Ising chain evolving under the transverse Ising Hamiltonian. We start from a chosen pure state $|\psi_{0}\rangle$ and one time step consists of evolution for time $\tau$ followed by the measurement. (Note that the measurement is performed with certainty at every time step, and the only source of stochasticity is the quantum nature of the system.) We study, as a function of time, three quantities, namely, bipartite entanglement, multipartite entanglement, and the survival probability of the initial state $|\psi_{0}\rangle$. From numerical studies on chains of size $L\leq 28$, it is found that (i) the bipartite entanglement shows a volume-to-area law transition at some critical value $\tau=\tau_{c}$ and (ii) the multipartite entanglement as well as the survival probability also show non-analytic behavior at the same value of $\tau$. Thus, we observe a ‘measurement-induced phase transition’ just like random circuit models Skinner et al. (2019); Koh et al. (2023). This transition point exists for every value of $h$ and is hence unrelated to the order-disorder transition at $h=1$. We present below the results for $h=\frac{1}{2}$ and include the results for $h=\frac{3}{2}$ in the appendix A. The central result of this work is that, exploiting the analytic nature of the Hamiltonian, we can calculate the survival probability at large $L$($\sim 1000$) too and show that the critical value $\tau_{c}$ scales with system size as $1/\sqrt{L}$. This proves that $\tau_{c}$ vanishes at the thermodynamic limit, and the transition we observe at small size exists for finite sizes only. Hence, whether the transition in random circuit models survives in the thermodynamic limit is an important question.
In the next section II, we present the measurement protocol. In Sections III and IV we present the analysis of bipartite entanglement and generalized geometric measure in small systems. Sec. V is devoted to the analytic theory, and Sec. VI is devoted to the summary.
II protocol and survival probability
Figure 1: $R_{n}$ vs $n$: Survival probability for different $\tau$ values : Here, we have plotted $R_{n}$ vs. $n$ graph corresponding to each $\tau$ value for system size $L=28$ at field $h=\frac{1}{2}$ where $R_{n}$ is the survival probability (Eq. 7) and $n$ is the number of cycle of the protocol.
We consider (at zero temperature) a chain of $L$ Ising spins in a (pure) state where all the spins are in the +Z direction:
$|\psi_{0}\rangle=|000\cdots 0\rangle$ (1)
It is allowed to evolve unitarily for a time $\tau$ under the transverse Ising Hamiltonian.
$\mathcal{H}=-J\sum_{j=1}^{L}s^{x}_{j}s^{x}_{j+1}-\Gamma\sum_{j=1}^{L}s^{z}_{j}$ (2)
(where $J$ is the interaction strength and $\Gamma$ is the strength of the transverse field. We fix $J=1$ and use symbol $h$ for $\Gamma/J$ ).
We now perform a projective measurement of the transverse magnetic moment per site $M_{z}=(1/L)\sum_{i=1}^{L}s_{i}^{z}$ and ask the question
Figure 2: Behavior of height of the plateau region of $R_{n}$ with $\tau$: In panel (a), we have plotted the $H$ vs. $\tau$ graph, where $H$ is denoted as the height of the plateau region for a given $\tau$ that is already depicted in FIG. 1. When there are multiple plateaus, we consider the height of the first one. Here, the black color denotes the curve corresponding to field $h=0.5$, and the orange color denotes the curve corresponding to $h=\frac{3}{2}$. The transition occurs around $\tau_{c}=0.2$. The point where this transition occurs is perfectly realized from the curve of panel (b), where we have plotted $\frac{dH}{d\tau}$ vs. $\tau$ for various system sizes, that is, $L=20,22,24,26$ at field $h=\frac{1}{2}$. Here, the highest value to each curve corresponds to the transition point.
Is the magnetic moment (per site) $M_{z}=1$? (3)
We perform the experiment on a large number ${\mathcal{N}}$ of identically prepared replicas, then note down the fraction of cases where the answer is yes and call this fraction $p_{1}$. We discard the replicas that gave answer yes and with the other replicas carry on the unitary evolution under ${\mathcal{H}}$ for another time interval $\tau$ (in units of $J/\hbar$). Then we perform the same measurement and note down the number of replicas, say, $p_{2}{\mathcal{N}}$, which yield yes. We repeat this procedure of unitary evolution, followed by projective measurement, so that after $n$ such cycles we get the number of replicas, say, $p_{n}{\mathcal{N}}$, which yield yes. Then, the probability that the answer is yes for the first time at the $n$-th measurement is $p_{n}$, and the probability that the system has yielded no at the first $n$ successive measurements is
$R_{n}\equiv 1-\sum_{k=1}^{n}p_{k}.$ (4)
We shall call $R_{n}$ the survival probability. The above protocol for dynamics has been used earlier Dhar et al. (2015a, b); Dhar and Dasgupta (2016) as a reasonable protocol for measuring the probability of first occurrence within the framework of quantum measurement.
Figure 3: $\tau_{c}$ vs $L$ and $h$ : Here, the critical point $\tau_{c}$ as we obtain from the survival probability $R_{n}$ is plotted with respect to the system size $L$ and field strength $h$.
Analysis for small systems: After the time evolution from $t=0$ to $t=\tau$, the system attains a state $\exp(-i\mathcal{H}\tau)|\psi_{0}\rangle$. Since $|\psi_{0}\rangle$ is the only state with $M_{z}=1$, the effect of the measurement stated in (3) is to project out $|\psi_{0}\rangle$ and get the state
$|\psi_{1}\rangle=e^{-i\mathcal{H}\tau}|\psi_{0}\rangle-\langle\psi_{0}|e^{-i\mathcal{H}\tau}|\psi_{0}\rangle|\psi_{0}\rangle$ (5)
at the completion of the first time step. Now the state $|\psi_{1}\rangle$ evolves again for time $\tau$ under $\mathcal{H}$, undergoes measurement, and becomes
$|\psi_{2}\rangle=e^{-i\mathcal{H}\tau}|\psi_{1}\rangle-\langle\psi_{0}|e^{-i\mathcal{H}\tau}|\psi_{1}\rangle|\psi_{0}\rangle$ (6)
The state function $|\psi_{n}\rangle$ after $n$ states is calculated numerically using the Chebyshev polynomial technique Piotr et al. (2019). and the survival probability is obtained therefrom as
$R_{n}=\langle\psi_{n}|\psi_{n}\rangle$ (7)
The plot of $R_{n}$ as a function of $n$ (for a given $\tau$) shows a plateau region at small $n$. As we increase the value of $\tau$, this region increases in value and saturates near 1 for a high value of $\tau$ (FIG. 1).
The principal quantity of our study is the rate at which the height of the plateau region varies with $\tau$ (FIG. 2(a)), and we observe that this quantity shows a peak at some value of $\tau$ (FIG. 2(b)).
Figure 4: Survival probability $R_{n}$ vs time-step $n$ : Here, the survival probability $R_{n}$ is plotted with respect to time-step $n$ for $h=\frac{3}{2}$. System size is $L=50$. A logarithmic decay at large time-steps is observed, which is not present at $h<1$.
When there are multiple plateaus, we consider the height of the first one. We note that the height of the peak rises with system size and identify this value of $\tau$ as the critical value $\tau_{c}$ (FIG. 2(b)) where the plot of $\frac{dH}{d\tau}$ vs $\tau$ is given. Throughout the paper, all the derivatives are calculated numerically by the two-point formula. We note that at finite size we have different $\tau_{c}$ at $h=\frac{1}{2}$ and $h=\frac{3}{2}$. The dependence of critical point $\tau_{c}$ on system size and $h$ is studied in FIG (3). It is observed that the dependence of $\tau_{c}$ on $h$ decreases as the system size $L$ increases. $\tau_{c}$ also moves towards smaller values as system size $L$ increases.
In FIG (4), the survival probability is studied at large $n$ at $h=\frac{3}{2}$ for different $\tau$ values. We can see that $R_{n}$ decays in logarithmic fashion at large $n$. This kind of logarithmic decay is not present in ferromagnetic phase ($h<1$) and can be observed only when the system is in paramagnetic phase ($h>1$). However, when $h$ is too high, such decay is not present. Thus, the pattern of decay can detect the different quantum phases of the ground state of transverse field Ising model.
III bipartite entanglement
Figure 5: Entanglement behavior with system size $L$ and $\tau$ : In panel (a), entanglement entropy for the whole system in the $L/4$ : rest bipartition ($S(L/4)$) is plotted for different $\tau$ values with respect to $L$, which is the system size. $\tau$ values increase as the color of the curves becomes deeper. We have taken $n=10$ for measuring $S$. It is apparent that there is an area law to volume law transition. In panel (b) we have plotted $S(L/4,\tau)$ for different values of $L$ with respect to $\tau$. In panel (c) we have shown the scaling collapse of the entanglement curves occurs when $\tau_{c}=0.2$ with appropriate exponents, which is $0.3$ following Ref. Skinner et al. (2019); Koh et al. (2023).
Bipartite entanglement refers to entanglement, a quantum correlation Horodecki et al. (2009); Gühne and Tóth (2009); Das et al. (2017) between two subsystems (or parties) in a composite quantum system. Suppose that $A$ and $B$ are two parts of a single quantum system, and that the state of the entire system is $\rho^{AB}$.The state is said to be entangled if it cannot be decomposed into $\rho^{AB}=\sum_{i}p_{i}\rho^{A}_{i}\otimes\rho^{B}_{i}$, for any ensemble $\{p_{i},\rho^{A}_{i}\otimes\rho^{B}_{i}\}$. For calculating the bipartite entanglement of a pure bipartite state, the measure we consider is the von Neumann entropy or entanglement entropy Bennett et al. (1996). Let us consider the two subsystems. $A$ and $B$ consist of $\ell$ and $L-\ell$ sites, respectively. Here, $L$ is the total number of sites. The measure, entanglement entropy, for any pure state of the Ising chain, after $n$ time-steps, is computed as
$S(n,\ell)=-\operatorname{Tr}[\rho_{A}\ln\rho_{A}]=-\operatorname{Tr}[\rho_{B}\ln\rho_{B}],$ (8)
where $\rho_{A}$ and $\rho_{B}$ are the reduced density matrices for the subsystems $A$ and $B$. Throughout the paper, we have taken log base e when we calculate bipartite entanglement. In this logarithmic base, the entanglement of bell states that is maximally entangled states is $0.693$ nats.
Following Srednicki (1993); Eisert et al. (2010), we know that the entanglement entropy of non-degenerate ground state of a gapped Hamiltonian, follow the area law Hastings (2007) (i.e., it saturates as the subsystem size increases), while a generic state of the system follows the volume law (i.e. changes linearly with subsystem size).
In the context of MIPT, a volume law to area law transition of the entanglement entropy occurs at a critical rate of measurement ($\mathcal{P}_{c}$) as defined in Ref. Skinner et al. (2019). According to Skinner et al. (2019), near the critical point, for certain set-ups, there exists a general scaling entanglement entropy, which can be written as $S(\mathcal{P})-S(\mathcal{P}_{c})=F\left((\mathcal{P}-\mathcal{P}_{c})L^{\frac{1}{\nu}}\right)$ where $L$ is the system size and $S$ is the entanglement entropy of the whole system in the $L/2:L/2$ bipartition, which is the von Neumann entropy of a subsystem of size $L/2$. The form of $F$ would depend on the set-up considered. The entanglement entropy depends more strongly on system size in the volume law phase, than in the area law phase. When $S(\mathcal{P})-S(\mathcal{P}_{c})$ is plotted against $(\mathcal{P}-\mathcal{P}_{c})L^{1/\nu}$, all the curves for different $L$ collapse on each other around $\mathcal{P}_{c}$, which indicates a phase transition at $\mathcal{P}_{c}$.
In this section, we have calculated the entanglement in the $L/4:\text{rest}$ biseparation of the entire system, and referred to it as $S(L/4,\tau)$, for different $\tau$ values Li et al. (2018). The entanglement is quantified by using the entanglement entropy, i.e., the von Neumann entropy of the $L/4$ subsystems, for different $L$. An area law to volume law transition can be observed when $S(L/4,\tau)$ is plotted with respect to $\tau$ (FIG. 5(a)). In FIG. 5(b), entanglements for different $L$ collapses to similar values at low $\tau$, and becomes system-size dependent at larger $\tau$, which also indicates an area law to volume law transition. In FIG. 5(c), $S(L/4,\tau)-S(L/4,\tau_{c})$ is plotted with respect to $(\tau-\tau_{c})L^{\mu}$. We find that scaling collapse occurs around $\tau_{c}=0.2$ and $\mu=0.3$, which shows that an entanglement area law phase transition point exists around $\tau_{c}=0.2$ (FIG. 5).
IV GENERALIZED GEOMETRIC MEASURE
For quantifying multipartite entanglement, we utilize the measure referred to as the generalized geometric entanglement (GGM) Sen (De); Das et al. (2016). See also Barnum and Linden (2001); Plenio and V. (2001); Wei and Goldbart (2003); Osterloh and Siewert (2005); Yu and Song (2006); Orús (2008). A pure multiparty quantum state is said to be genuinely multiparty entangled if it cannot be factorized into a product state in any biseparation. The GGM of any pure multiparty state is defined as the minimum distance of the given state from the set of all states that are not genuinely multiparty entangled, with the Bures metric being used to obtain the distance function. We are interested in analyzing the genuine multiparty entanglement of the pure $L$-partite state, $|\psi_{n}\rangle$, after the $n$th measurement.
The GGM of |$\ket{\psi_{n}}$ is defined as
$G(n)=1-\Lambda^{2}_{max}(\ket{\psi_{n}}),$ (9)
where $\Lambda_{max}(\ket{\psi_{n}})$ is the maximum fidelity between the given state |$\ket{\psi_{n}}$ and an arbitrary non-genuinely multiparty entangled state |$\ket{\xi}$. So, a maximization is to be performed over all non-genuinely multiparty entangled pure states, which, however, can be explicitly obtained in closed form Sen (De):
$G(n)=1-\max_{\ell}\left\{\lambda_{\ell,L-\ell}^{2}\right\}_{l}$ (10)
where $\lambda_{\ell,L-\ell}$ is the maximal Schmidt coefficient in the bipartite split $\ell:L-\ell$ of |$\ket{\psi_{n}}$.
Since we calculate the GGM for the state $|\psi_{n}\rangle$ after the $n$th measurement, we plot the “stochastically attained GGM” (SAG), defined as
$S_{G}(n)=R_{n}\times G(n)$
The plot of $S_{G}(n)$ with $n$ at field strength $h=1/2$ is depicted in FIG. 6(a) for different $\tau$ values, viz. $\tau=0.1,0.2,0.3,\ldots,1.0.$ A similar trend, that is, a large gap between the curves at $\tau$ values $0.1$ and $0.2$, is noticed, like the $R_{n}$ vs $n$ graph in FIG. 1. To further analyze the behavior of multipartite entanglement, we perform cumulative averaging, which, for a sequence, $\bigl{\{}x_{1},x_{2},x_{3},\ldots,x_{n}\bigl{\}}$, is defined as $\mathcal{M}_{i}=\frac{1}{i}\sum_{k=1}^{i}x_{k}$. Therefore, $\mathcal{M}_{i}$ is the cumulative average at the $i$th step.
In our case, the cumulative average of $S_{G}(n)$ is denoted as $\mathcal{G}_{n}$, and in FIG. 6(b), we have plotted the $\mathcal{G}_{n}$ against $n$ for the same set of $\tau$ values as in FIG. 6(a) for the same field strength, that is $h=\frac{1}{2}$. The gap between $\tau=0.1$ and $\tau=0.2$ curves becomes more prominent after we consider the cumulative values. For $\tau=0.2$ and onwards, the gap starts to decrease, and from $\tau=0.3$, all the curves overlap on each other. We point out that the curves show a plateau region, just like survival probability for small $n$ (FIG. 1). This feature of the curves, corresponding to different values of $\tau$, leads us to check the behavior of $\frac{d\mathcal{G}}{d\tau}$ with $\tau$. We identify the value of $\tau$ corresponding to the highest value of the derivative $\frac{d\mathcal{G}}{d\tau}$ as $\tau_{c}$.
Thus, for sizes $L\leq 28$, the three quantities, namely survival probability, bipartite entanglement entropy, and genuine multipartite entanglement, show a critical behavior at a common value of $\tau_{c}$ around $0.2$ at field strength $h=1/2$, and this we consider as the signature of a measurement-induced transition in this quantum many-body system. These results resemble those in the studies on random circuit systems with local stochastic measurements Koh et al. (2023); Skinner et al. (2019). We have repeated the entire analysis for all the three mentioned order parameters at field strength, $h=\frac{3}{2}$, and the results are elaborated in Appendix A.
Figure 6: Behavior of multipartite entanglement with time-step $n$ and $\tau$ : In panel (a), we have plotted $S_{G}(n)$ vs $n$ at field strength $h=\frac{1}{2}$ for different $\tau$ values. Here, the $\tau$ values are taken from $\tau=0.1,0.2,.....,1.0$. A gap between the curves corresponding to the $\tau$ values is detected. Next, in panel (b), $\mathcal{G}_{n}$ vs $n$ is plotted. Here, $\mathcal{G}_{n}$ is the cumulative average of the sequence $\bigl{\{}S_{G}(1),S_{G}(2),S_{G}(3),S_{G}(4)...S_{G}(n)\bigl{\}}$. In panel (c) we have plotted $\frac{d\mathcal{G}_{n}}{d\tau}$ vs $n$ plot, where the peak value occurs around $\tau=\tau_{c}=0.2$. It depicts the transition that happens near $\tau_{c}=0.2$.
V Recursion relation for survival probability
It has been shown earlier Dhar and Dasgupta (2016) that the transition when the measurement is complementary to the question (3), namely,
Is the magnetic moment (per site) $M_{z}\neq 1$? (11)
can be treated analytically for any arbitrary system size. It was proved that the ratio of two successive first detection probabilities is given by
$\frac{p_{n}}{p_{n-1}}=e^{-\alpha L\tau^{2}}$ (12)
with
$\alpha(h,\tau)=-\frac{1}{L\tau^{2}}\sum_{k=0}^{\pi}\,\log\left[1-\left(\frac{2\sin k\,\sin(\lambda_{k}\tau)}{\lambda_{k}}\right)^{2}\right]$ (13)
The quantity $\alpha$ shows a change of slope, indicating a criticality, as a function of $h$ and $\tau$. However, as is evident from Eq. (12) the first detection probability is finite only for finite size and vanishes in the thermodynamic limit.
For the present measurement (3), we could not derive a closed-form expression for the survival probability, but we could derive a recursion relation that enables one to compute the probability $R_{n}$ for large size. To this end, we introduce the quantities
$|\phi_{n}\rangle=e^{-i\mathcal{H}\tau n}|I\rangle,\;f_{n}=\langle I|e^{-i\mathcal{H}\tau n}|I\rangle,\;n=0,1,2,\cdots$ (14)
Then, $|\phi_{0}\rangle=|I\rangle$, $f_{n}=\langle I|\phi_{n}\rangle$, $\langle\phi_{m}|\phi_{n}\rangle=f_{n-m}$ and
$|\psi_{1}\rangle=|\phi_{1}\rangle-f_{1}|I\rangle$ (15)
After the second time-step,
$|\psi_{2}\rangle$ $=$ $e^{-i\mathcal{H}\tau}|\psi_{1}\rangle-\langle I|e^{-i\mathcal{H}\tau}|\psi_{1}\rangle|I\rangle$ (16)
$=$ $|\phi_{2}\rangle-f_{1}|\phi_{1}\rangle+(f_{1}^{2}-f_{2})|I\rangle$
The basic idea is to observe that for $n=0,1,2$, the wave-function can be expressed as
$|\psi_{n}\rangle=\sum_{m=0}^{n}C_{m}^{(n)}|\phi_{m}\rangle$ (17)
and that the next wave-function
$|\psi_{n+1}\rangle=e^{-i\mathcal{H}\tau}|\psi_{n}\rangle-\langle I|e^{-i\mathcal{H}\tau}|\psi_{n}\rangle|I\rangle$ (18)
can also be written as,
$|\psi_{n+1}\rangle=\sum_{m=0}^{n+1}C_{m}^{(n+1)}|\phi_{m}\rangle$ (19)
with the recursion relations
$C^{(n+1)}_{0}=-\sum_{m=0}^{n}C^{(n)}_{m}f_{m+1},\;\;\;C^{(n+1)}_{m}=C^{(n)}_{m-1}$ (20)
for $0<m\leq n$. Note that
$C_{n}^{(n)}$ $=$ $C^{(n+1)}_{n+1}=1,$ (21)
$C_{0}^{(1)}$ $=$ $-f_{1},\;\;\;C_{1}^{(2)}=-f_{1},\;\;\;C_{0}^{(2)}=f_{1}^{2}-f_{2}$ (22)
and that
$C_{m}^{(n)}=C_{m-1}^{(n-1)}=C_{0}^{(n-m)}$ (23)
Using the recursion relations, one can find the survival probability after $n$ measurements as
Figure 7: Survival probability at large system sizes : In panel (a), we have plotted the derivative of $H$ vs. $\tau$ graph for $L=100,200,300,500,1000$, where $H$ is denoted as the height of the plateau region for a given $\tau$ obtained by using Eq. (24). It is apparent that $\tau_{c}$ moves towards 00 as $L$ increases. In panel (b), we have plotted $\frac{dH}{d\sigma}$ vs. $\sigma$ for various system sizes, that is, $L=100,200,300,500,1000$ at field $h=\frac{1}{2}$ where $\sigma=\tau\sqrt{L}$. All the curves have a peak around $\sigma_{c}=1$
$R_{n}=\langle\psi_{n}|\psi_{n}\rangle=\sum_{m_{1},m_{2}=0}^{(n)}\left(C_{m_{1}}^{(n)}\right)^{*}C_{m_{2}}^{(n)}\,f_{m_{2}-m_{1}}$ (24)
This expression is valid for any Hamiltonian, provided $f_{n}$ are known. For the specific case of transverse Ising Hamiltonian, one can derive an expression for $f_{n}$ in closed form for any even value of $L$, by using the exact solution Damski and Rams (2013). One can transform spin variables $s_{j}$ to fermion variables $a_{j}$ and perform Fourier transformation to fermion variables $a_{k}$, to get $\mathcal{H}$ as a Kronecker sum of commuting Hamiltonians $\mathcal{H}_{k}$ :
$\mathcal{H}$ $=$ $\sum_{k=0}^{\pi}\mathcal{H}_{k},\;\;\mathcal{H}_{k}=-2i\sin k\left[a_{k}^{\dagger}a_{-k}^{\dagger}+a_{k}a_{-k}\right]$ (25)
$-2(\Gamma+\cos k)\left[a_{k}^{\dagger}a_{k}+a_{-k}^{\dagger}a_{-k}-1\right]$
where $k=(2n+1)\pi/L$ with $n=0,1,2,\cdots,L/2-1$. Each of these $\mathcal{H}_{k}$’s can be described by four basis states namely, $|00\rangle_{k}$, $|11\rangle_{k}$, $|10\rangle_{k}$ and $|01\rangle_{k}$, where the numbers in each basis signify the occupation status of the fermions having momenta $+k$ and $-k$ respectively. (We consider the even-occupation states only.) The ground state and excited state of $\mathcal{H}_{k}$ are
$|GS\rangle_{k}$ $=$ $i\cos\theta_{k}|11\rangle_{k}-\sin\theta_{k}|00\rangle_{k},$
$|ES\rangle_{k}$ $=$ $i\sin\theta_{k}|11\rangle_{k}+\cos\theta_{k}|00\rangle_{k}$ (26)
[…]
$e^{-i\mathcal{H}_{k}n\tau}|11\rangle_{k}$ $=$ $\left[\cos(\lambda_{k}n\tau)+i\sin(\lambda_{k}n\tau)\cos(2\theta_{k})\right]|11\rangle_{k}$ (27)
$-\sin(2\theta_{k})\sin(\lambda_{k}\tau)|00\rangle_{k}$
This gives finally,
$f_{n}=\prod_{k}\left[\cos(\lambda_{k}n\tau)+i\sin(\lambda_{k}n\tau)\cos(2\theta_{k})\right]$ (28)
In principle, one should be able to calculate the survival probability using Eqs. (20, 24, 28), but due to some precision problems we could not go beyond $L=1000$ and $n=10000$.
As for small size, the survival probability shows a plateau region and the derivative of the height $H$ of the plateau with respect to $\tau$ shows a peak at a critical point $\tau=\tau_{c}$. (Fig. 7). It is found that if one uses instead of $\tau$ a scaled variable, $\sigma=\tau\sqrt{L}$ then the plot of $dH/d\sigma$ vs $\sigma$ coincide for different values of $L$. This indicates that the position of the peaks $\tau_{c}$ varies as $1/\sqrt{L}$ and vanishes for large $L$. We shall derive below this scaling law in the limit of small $\tau$ and small $n\tau$.
To this end, we first write $f_{n}$ in Eq. (28) as
$f_{n}=\rho_{n}e^{i\Phi_{n}}$
and note that
$\rho_{n}=\exp\left[\frac{1}{2}\sum_{k}\log\left[1-\left(\frac{2\sin k\,\sin(\lambda_{k}n\tau)}{\lambda_{k}}\right)^{2}\right]\right]$ (29)
$\Phi_{n}=\sum_{k}\tan^{-1}\left[\frac{2(h+\cos k)}{\lambda_{k}}\,\tan(\lambda_{k}n\tau)\right]$ (30)
For small values of $\tau$ and $n$, one obtains
$\Phi_{n}=\sum_{k}2n\tau(h+\cos k)=\mu n$
where $\mu=2\tau hL$. The recursion relations then tell us,
${\rm Arg}\left(C^{(n)}_{m}\right)=\mu(n-m)$
and hence the argument of each summand in the expression for $R_{n}$ in Eq. (24) is zero. Thus, the main contribution to $R_{n}$ comes from $\rho_{n}$, which, in the limit of small $n\tau$ becomes
$\rho_{n}=\exp\left[-\frac{1}{2}n^{2}\tau^{2}L\right]$
This indicates that the parameters $\tau$ and $L$ may be expected to occur as the combination $\tau^{2}L$ in the value of survival probability.
VI Summary
We present the observation of measurement-induced phase transition in a quantum spin model with projective non-random measurement. We start with a chain of $L$ Ising spins all oriented in the +Z direction. In each time step, it evolves unitarily under the transverse Ising Hamiltonian for time $\tau$, and then a measurement is performed that gives a binary answer to the question, “Is the magnetic moment per site equal to 1?”. After $n$ time-steps, the probability that the answer is “no” in each case is termed as the survival probability $R_{n}$. We have derived a recursion relation, using which we could compute $R_{n}$ for size up to $L=1000$. The plot of $R_{n}$ vs. $n$ shows a plateau region at small $n$. The key parameter is the rate at which the height (say $H$) of this plateau changes with $\tau$. The plot of $\frac{dH}{d\tau}$ shows a peak at some value of $\tau$, which we identify as the critical point $\tau_{c}$, since the bipartite and multipartite entanglement undergoes an area-law to volume-law transition there. We prove analytically that, for small $\tau$, the critical point $\tau_{c}$ scales as $1/\sqrt{L}$. On using the scaled variable $\sigma=\tau\sqrt{L}$, one finds that the plots of $\frac{dH}{d\sigma}$ vs $\sigma$ for different system sizes are coincident, and as $L$ increases, the peak in the plot of $\frac{dH}{d\tau}$ rises in height and shifts towards zero. Hence, the transition is only observable for finite $L$. Interestingly, the survival probability decays logarithmically with time in some cases. All these cases correspond to the phase for which the ground state of the Hamiltonian is disordered, indicating that this transition is dependent on the presence of order in the ground state of the Hamiltonian. However, there are some values of the external field for which the ground state is disordered, but the logarithmic decay is not present. Also, one must note that all the results discussed above are for one particular choice of the initial state.
Acknowledgements
PN acknowledges UGC, India for financial support (Ref. No. 191620072523) and Harish Chandra Research Institute for access to their infrastructure.
Appendix A
Here, in appendix A, we do the same analysis as the main paper but at field strength $h=\frac{3}{2}$. Firstly, in FIG. A.1, we have shown the $R_{n}$ vs. $n$ plot at $h=\frac{3}{2}$. Here, the notation $R_{n}$ and $n$ are consistent with the main paper. The same feature of the curves corresponding to the $\tau$ value $0.1$ and $0.2$ is captured.
Figure A.1: $R_{n}$ vs $n$: Here, we have plotted $R_{n}$ vs. $n$ graph corresponding to each $\tau$ value for system size $L=28$ at field $h=\frac{3}{2}$.
Figure A.2: $\frac{dH}{d\tau}$ vs $\tau$ plot. We have plotted $\frac{dH}{d\tau}$ vs. $\tau$ for different system sizes, that is, $L=20,22,24,26,28$ at field $h=\frac{3}{2}$. The curves precisely illustrates the $\tau$ values at which this transition happens. The transition point in this case is represented by the highest values of the curves.
We again plot the derivative of the height of the plateaus from FIG. A.1 with respect to $\tau$. Then, we see the behavior of $\frac{dH}{d\tau}$ vs $\tau$ in FIG. A.2 for various system sizes, that is, $L=20,22,24,26,28$. Here, the peaks of the curves are identified as the critical point. The transition point moves towards a lower value of $\tau$ as it is seen in the study for $h=\frac{1}{2}$.
Figure A.3: In panel (a), entanglement entropy for the whole system in $L/4:\text{rest}$ bipartition ($S(L/4)$) is plotted for different $\tau$ values with respect to $L$, which is the system size. We have taken $n=10$ and calculated entanglement and see an area law to volume law transition. $\tau$ values increase as the color of the curve becomes deeper. In panel (b) we have plotted $S(L/4,\tau)$ for different values of $L$ with respect to $\tau$. In panel (c) we have shown the scaling collapse of the entanglement curves occurs around $\tau_{c}=0.1$ with the exponent $0.4$.
Figure A.4: In panel (a), the plot of $S_{G}(n)$ vs $n$ at field strength $h=\frac{3}{2}$ for different $\tau$ values. Here, the $\tau$ values are taken from $\tau=0.1,0.2,.....,1.0$. A gap between the curves corresponding to the $\tau=0.1$ and $\tau=0.2$ values is detected. Next in panel (b), $\mathcal{G}_{n}$ vs $n$ is plotted at the same field strength, that is, $h=\frac{3}{2}$. In panel (c) we have plotted $\frac{d\mathcal{G}_{n}}{d\tau}$ vs $n$ for $L=26$. plot, which shows the transition point around $\tau_{c}=0.1$.
After that we calculate bipartite entanglement and show from FIG. 3(a) the area to volume law transition occurs apparently at $\tau_{c}=0.1$. at field strength $h=\frac{3}{2}$. This area-to-volume law transition is also detected from the collapse of curves of different system sizes $L$ at $\tau<\tau_{c}=0.1.$. It is shown at 3(b). At last in FIG. 3(c) we show the scaling collapse at $\tau_{c}=0.1$ and $\mu=0.4$.
Then we also analyze stochastically attained GGM (SAG) at the field strength $h=\frac{3}{2}$. In FIG. 4(a) we have plotted $S_{G}(n)$ vs $n$ for ten $\tau$ values from $0.1$ to $1.0$ with an interval $0.1$. and in FIG. 4(b)  $\mathcal{G}_{n}$ vs $n$ plot is depicted for the same set of $\tau$ values as shown in FIG. 4(a). Here also, the notations $S_{G}(n)$ and $\mathcal{G}_{n}$ are consistent with the main text. At last in FIG. 4(c) from $\frac{d\mathcal{G}_{n}}{d\tau}$ curves for system size $L=26$, the transition occurs around the point $\tau_{c}=0.1$.
Figure A.5: In panel (a), we have plotted $\frac{dH}{d\tau}$ vs. $\tau$ for $L=100,200,300,500,1000$ at the field strength $h=\frac{3}{2}$. Even at $h=\frac{3}{2}$, $\tau_{c}$ also moves towards 00 as $L$ increases. In panel (b), we have plotted $\frac{dH}{d\sigma}$ vs. $\sigma$ at the same system sizes and field strength as FIG. 5(a), where $\sigma=\tau\sqrt{L}$. All the curves have a peak around $\sigma_{c}=1$.
At last, in FIG. A.5 we do some large system size studies at field strength at $h=\frac{3}{2}$. In FIG. 5(a) we show the variation of $\frac{dH}{d\tau}$ with $\tau$ at field strength $h=\frac{3}{2}$ for $L=100,200,300,500,1000$. The highest value of $\frac{dH}{d\tau}$ vs $\tau$ curve shows the deviation of the transition point $\tau_{c}$ towards 00 with increase of the system size. In FIG. 5(b) we show the $\frac{dH}{d\sigma}$ vs. $\sigma$ plot. Here the quantity $\sigma=\tau\sqrt{L}$. As all the curves corresponding to different system sizes have peak values at $\sigma_{c}=1$, it shows that $\tau_{c}\sim\frac{1}{\sqrt{L}}$.